# ApisTox: a new benchmark dataset for the classification of small molecules toxicity on honey bees

### v1

@@ Methods
ApisTox dataset is available on GitHub [25], as well as on Zenodo [26]. We distribute the files on all stages of processing: raw, cleaned, the final dataset, and data splits as described in Methods section. All files are in human-readable CSV format. Main dataset file is dataset_final.csv, in outputs directory, with structure described in Table 1.
Raw files are in raw_data directory. ecotox.csv contains raw outputs of ECOTOX database query, and uses pipe “|” as separator. All other files use commas as separators. bpdb.csv and ppdb.csv consist of data from BPDB and PPDB, respectively.
Processed data files are in outputs directory. ecotox_cleaned_data.csv is ECOTOX data after processing and cleaning, and excluded_data.csv contains rows removed at any point of processing, along with reason for exclusion. dataset_final.csv is the main dataset file.
@@
For each split, described in Methods section, we provide training and testing subset of data after splitting in outputs/splits directory. All those files have exactly the same structure as the main dataset file.
@@
In this section, we present molecular characteristics of ApisTox dataset, and technical analyses with basic properties relevant to chemoinformatical applications. We use the data from dataset_final.csv file. Suggestions for further modelling and applications are in the Usage Notes section.
@@
Here, we perform the basic quality checks for molecular data. ApisTox, as a curated and unified collection of data, should cover all three source databases. We compare the number of molecules of the final dataset and source databases, all after the same data cleaning procedure outlined in the Methods section, in Table 2. ApisTox is indeed larger than all input datasets, validating our data combination process. In particular, it is also almost 25% larger than PPDB, the largest of the source datasets.
@@
For more detailed analysis, we verify that all canonical SMILES from those cleaned source datasets are included in the final dataset. Results are summarized by a Venn diagram in Figure 2. We gain molecules from each data source, showing that creation of unified datasets using many input databases is beneficial for ecotoxicology data. Almost molecules are included in the final dataset, except for 11 SMILES strings from ECOTOX. However, this is not a mistake - manual check showed that those molecules are subtle duplicates, having the same CAS numbers as molecules from PPDB already included. This is the consequence of non-uniqueness of SMILES format, i.e. a given molecule can be written in many ways, and RDKit canonicalization is not always able to detect such cases. This shows that the two-step deduplication, described in the Methods section, which includes both canonical SMILES and CAS deduplication, is indeed required for proper merging of molecular datasets.
Figure 2: Number of common molecules between source datasets and ApisTox.
We also verify that ApisTox contains only valid entries that can be processed by RDKit, and contains no duplicated molecules (in terms of canonical SMILES and CAS numbers). To further validate our preprocessing workflow in this regard, we apply it to datasets previously proposed for honey bee toxicity: CropCSM [3], BeeTOX [1] and BeeToxAI [7]. As shown in Table 3, other datasets contain invalid SMILES and duplicates, even as much as 36% for BeeToxAI. Furthermore, ApisTox is considerably larger than all of them, containing many more toxic molecules in particular, which are crucial for understanding the underlying causes of pesticide toxicity for honey bees.
@@
We present distributions of toxicity labels in Figure 4. In terms of binary toxicity label, ApisTox is moderately imbalanced, with pesticides toxic for honey bees constituting 29% of the data. Concretely, there are 739 non-toxic and 296 toxic compounds. The ternary level, following PPDB methodology, is more severely imbalanced, with 17% non-toxic, 66% moderately toxic, and 17% highly toxic molecules. This is due to the very wide definition of moderate toxicity in PPDB. Under this methodology, we have 177 non-toxic, 687 moderately toxic and 171 highly toxic molecules.
Such imbalance in class distributions influences metrics appropriate for validating predictive models, as discussed in the Usage Notes section.
@@
We analyze the effects of different proposed splits into training and testing sets. Preferably, both parts should have similar distribution of classes and pesticide types, to keep the basic characteristics of the dataset similar and have good representation of all segments of the data in the test set.
@@
Class distributions for both binary labels and ternary PPDB levels are similar in all cases, which indicates that no splitting method introduces unwanted additional imbalance. The diversity is highest for MaxMin split and the lowest for stratified random split, which follows their motivation outlined in the Methods section. Train-test separation is also the highest for MaxMin split, meaning that not only its test set covers the chemical space of the dataset well, but also at the same time is not too similar to training compounds.
@@
Here we present analyses concerning internal structure of the molecules, i.e. their Bemis-Murcko scaffolds, functional groups and frequent subgraphs. High quality molecular datasets should, in general, be structurally valid and not be dominated by a few substructures. At the same time, we expect the existence of discriminative structures common only among pesticides toxic or non-toxic for honey bees, which would enable data analysis and interpretability of predictive algorithms.
Firstly, we validate the distribution of molecular scaffolds, which motivated our omission of scaffold split in the Methods section. Among 1035 molecules, there are 424 Bemis-Murcko scaffolds, and 324 occur in only one molecule, indicating a very diverse dataset. This is also quite high number compared to typical medicinal chemistry datasets. Within the group of toxic molecules, scaffolds composed of 6-membered aromatic rings containing carbons and nitrogens (e.g.“c1ccccc1”, “c1cncncnc1”, “c1ccncc1”, “c1ncncn1”) are common, constituting 22% (67 out of 296) of these molecules.
However, Bemis-Murcko scaffolds are defined only for connected compounds with ring systems, and 186 molecules (almost 20% of the ApisTox dataset) have no scaffold at all (141 have multiple fragments, 45 do not have any rings). In particular, it means that the ones with disconnected components could share an entire subgraph with those with scaffolds available, and the scaffold split method would not be able to detect that.
Next, we check the distribution of the functional groups, also called fragments in RDKit. In Figure 6, we visualize ten most discriminative functional groups, i.e. those with the largest frequency difference between pesticides toxic and non-toxic to honey bees.
We see that there is a high proportion of insecticide-specific functional groups in toxic molecules. Phosphorus is very commonly found in instecticides, most often in a form of organophosphates [35, 36] (phos_ester and phos_acid fragments). Rich presence of sulfide fragment is explained by common usage of sulphur in pesticides, in particular fungicides [37, 38]. Pyrethrins and pyrethroids constitute a large group of insecticides [39], to which bees are very sensitive [40], containing an ester group and also very often a cyano group (ester and nitrile fragments). Neonicotinoids are yet another large group of insecticides, very rich in nitrogen, within which a hydrazine group is often detected, as well as a guanidinium group [41, 42] (hdrzine and guanido fragments). All of those results show that ApisTox data aligns with literature on pesticide toxicity of honey bees.
@@
Lastly, we apply molecular frequent subgraph mining, in order to identify discriminative substructures, i.e. both frequent overall and at the same time much more frequent among one toxic molecules than non-toxic ones (or vice versa). This is a more data-driven solution compared to functional groups, since we derive subgraphs from the data itself. It can also detect smaller subgraphs compared to typical functional groups. Again, we expect the data to reflect the usage of common chemical elements and their influence on pesticide toxicity to honey bees.
We use the MoSS tool [43] with default settings, to mine the most common subgraphs. We then identify their frequency in toxic and non-toxic classes, and select ten with the largest frequency difference between classes. Results are presented in Figure 7. The results again align with literature, with phosphate-containing subgraphs much more common among toxic compounds, and sulfur often found in both types of compounds.
@@
The important application of ApisTox dataset is understanding and predicting agrochemical compounds’ toxicity for honey bees, using data mining and machine learning (ML) methods. Those tasks can be broken into unsupervised and supervised applications, depending on whether they explicitly model the dependence between molecule features and selected targets or not.
@@
Analytical tasks include detecting functional groups, scaffolds, binding sites, and other molecular fragments that influence the toxicity of chemicals for honey bees. Those analyses can be carried out either on the entirety of the data, or on its subsets, e.g. for herbicides, fungicides, or insecticides only. Understanding those factors can lead to design of safer pesticides in the future. Techniques like graph clustering, dimensionality reduction, and frequent subgraph mining can be utilized [44, 45]. We stress that those methods, even when used as unsupervised learning techniques, often have hyperparameters and settings to tune, and they should also be validated on external data not used during the initial analysis. Therefore, provided train-test splits still have to be used, even if explicit labels are not utilized. In this regard, time split would be especially useful, as a direct approximation of the process of developing new agrochemicals.
@@
Predictive models for modelling toxicity of pesticides will most likely use label or ppdb_level columns as targets. They should utilize only the training data for parameter estimation. For hyperparameter tuning, cross-validation with stratified sampling is recommended, due to class imbalance and relatively small dataset size. Test data must not be used at any point until the final validation of generalization performance. We recommend using MaxMin split for this purpose, since it covers the chemical space of the dataset more uniformly than other splits, requiring good generalization across the entire domain of the dataset [20, 21].
@@
For measuring the performance of predictive models, metrics that take into consideration class imbalance should be used, and at least two or three metrics should be reported to take into consideration different aspects of model performance. In particular, we recommend Area Under Receiver Operating Characteristic curve (AUROC), since it works well with imbalanced data and also takes into consideration probabilistic outputs of classifiers, as well as Matthews Correlation Coefficient (MCC), which can sometimes detect model failures despite high AUROC value [46, 47]. Among other popular metrics, F1-score, precision, and recall can be used.
@@
ApisTox can also be used as a benchmark dataset for molecular graph classification. Since datasets and benchmarks in this area come almost exclusively from medicinal chemistry, performance of many models, like molecular fingerprints, graph kernels, and graph neural networks (GNNs), has been evaluated exclusively on tasks directly related to pharmacological de novo drug design. Our dataset allows validation of the generalization performance of such models on new domains of agrochemistry and ecotoxicology. In this context, usage of unified and predefined split and metrics is of paramount importance, to allow comparison of different models. We recommend usage of MaxMin split, as well as reporting both AUROC and MCC.
@@
When using ppdb_level as target variable, we stress that this is not a three class classification problem, but rather an ordinal regression problem, also known as ordinal classification [48]. Toxicity levels are ordered integers, and classes 0 and 2 (non-toxic and highly toxic) are more distant than 1 and 2 (moderately toxic and highly toxic). Therefore, appropriate models should be used, e.g. ordinal logit model instead of logistic regression. In this case, additionally reporting regression metrics like MAE (Mean Absolute Error) and RMSE (Root Mean Squared Error) is recommended, with additional corrections for class imbalance [49].
@@
Code is available on GitHub at https://github.com/j-adamczyk/apis_tox_dataset. Code uses Python 3.10. To ensure full reproducibility, we pinned all external library dependencies (including transitive dependencies) using Poetry dependency manager [50]. We also include poetry.lock file with all dependency versions, as well as requirements.txt file exported from it.
@@ Acknowledgements
Research was supported by the grant from “Excellence Initiative - Research University” (IDUB) for the AGH University of Krakow.
@@
J.A. and J.P.: Conceptualization, Investigation, Methodology, Data Curation, Software, Validation, Visualization, Writing - Original Draft, Writing - Review & Editing. P.S.: Supervision, Methodology, Writing - Original Draft, Writing - Review & Editing. All authors reviewed and approved the manuscript.

### v3

@@ Methods
ApisTox dataset is available on Zenodo [53], as well as on GitHub [54]. In both places, we distribute the files on all stages of processing: raw, cleaned, the final dataset, and data splits as described in the “Methods” section. All files are in human-readable CSV format. The main dataset file is dataset_final.csv (in outputs directory on GitHub), with columns described in Table 1.
ecotox.csv contains raw outputs of ECOTOX database query, and uses pipe “|” as separator. All other files use commas as separators. ppdb.csv and bpdb.csv files consist of data from PPDB and BPDB, respectively. For those two databases, we distribute only the subset of molecules that contained any bee pesticide data: 860 from PPDB and 122 from BPDB. From raw data available on the website, we only include the public identifiers: name, CID, CAS, SMILES. Columns “herbicide”, “fungicide”, “insecticide”, “other_agrochemical”, “label” and “ppdb_level” are features created from HTML responses, as described in the “Methods” section. On GitHub, those files are in the raw_data directory.
ecotox_cleaned_data.csv file contains ECOTOX data after processing and cleaning, and excluded_data.csv contains rows removed at any point of processing, along with reason for exclusion. dataset_final.csv is the main dataset file. On GitHub, those files are in the outputs directory.
For data splits, described in the “Methods” section, we provide training and testing subset of data after splitting, in files random_train.csv, random_test.csv, time_train.csv, and time_test.csv, maxmin_train.csv, maxmin_text.csv. Those files have exactly the same structure as the main dataset file. On GitHub, they are in outputs/splits directory.
@@
In this section, we present molecular characteristics of ApisTox dataset, and technical analyses of basic properties relevant to chemoinformatical applications. We use the data from dataset_final.csv file. Suggestions for further modeling and applications are in the “Usage Notes” section.
@@
Here, we perform the basic quality checks for molecular data. ApisTox, as a curated and unified collection of data, should cover all three source databases. We compare the number of molecules of the final dataset and source databases in Table 2, all after the same data cleaning procedure outlined in the “Methods” section. ApisTox is indeed larger than all input datasets, validating our data combination process. In particular, it is also almost 25% larger than PPDB, the largest of the source datasets.
@@
Furthermore, we verify that all canonical SMILES from the cleaned source datasets are included in the final dataset. Results are summarized by a Venn diagram in Figure 2. We integrated molecules from each data source, creating a significantly larger and unified dataset beneficial for ecotoxicology research. The final dataset comprises all filtered molecules from sourced databases, except for 11 SMILES strings from ECOTOX. Manual checks showed that those molecules are subtle duplicates, having the same CAS numbers as molecules from PPDB already included. This is the consequence of non-uniqueness of SMILES format, i.e. a given molecule can be written in many ways, and even RDKit standardization and canonicalization workflow is not always able to detect such cases. This shows that a two-step deduplication procedure, which includes both canonical SMILES and CAS deduplication, is indeed required for proper merging of molecular datasets (see the “Methods” section for details).
Figure 2: Number of molecules present in ApisTox sourced from three databases.
We also verify that ApisTox contains only valid entries that can be processed by RDKit, and contains no duplicated molecules (in terms of canonical SMILES and CAS numbers). To further validate our preprocessing workflow in this regard, we apply it to datasets proposed for honey bee toxicity analyses: CropCSM [5], BeeTOX [8] and BeeToxAI [19]. As shown in Table 3, other datasets contain invalid SMILES and duplicates, even as much as 36% for BeeToxAI. This emphasizes the need for highly rigorous data processing workflows like ours. ApisTox is considerably larger than all of them, containing many more toxic molecules in particular, which are crucial for understanding the underlying causes of pesticide toxicity for honey bees.
@@
We present distributions of toxicity labels in Figure 4. In terms of binary toxicity label, ApisTox is moderately imbalanced, with pesticides toxic for honey bees constituting 29% of the data. Concretely, there are 739 non-toxic and 296 toxic compounds. The ternary level is more severely imbalanced, with 17% non-toxic, 66% moderately toxic, and 17% highly toxic molecules. This is due to the very wide definition of moderate toxicity in PPDB. Under this methodology, we have 177 non-toxic, 687 moderately toxic and 171 highly toxic molecules.
Such imbalance in class distributions influences metrics appropriate for validating predictive models, as discussed in the “Usage Notes” section.
@@
We analyze the effects of different proposed splits into training and testing sets. Preferably, both parts should have similar distribution of classes and pesticide types, to keep the basic characteristics of the dataset similar and have good representation of all data segments in the test set.
@@
Class distributions for both binary labels and ternary PPDB levels are similar in all cases, which indicates that no splitting method introduces unwanted additional imbalance. The diversity is highest for MaxMin split and the lowest for stratified random split, which follows their motivation outlined in the “Methods” section. Train-test separation is also the highest for MaxMin split, meaning that not only its test set covers the chemical space of the dataset well, but also at the same time is not too similar to training compounds.
@@
Here, we present analyses concerning internal structure of the molecules, i.e. their Bemis-Murcko scaffolds, functional groups and frequent subgraphs. High quality molecular datasets should, in general, be structurally valid and not be dominated by a few substructures. At the same time, we expect the existence of discriminative structures common only among pesticides toxic or non-toxic for honey bees, which would enable data analysis and interpretability of predictive algorithms.
Firstly, we validate the distribution of molecular scaffolds, which motivated our omission of scaffold split in the “Methods” section. Among 1035 molecules, there are 424 Bemis-Murcko scaffolds, and 324 occur in only one molecule, indicating a very diverse dataset. This is also quite high number compared to typical medicinal chemistry datasets. Within the group of toxic molecules, scaffolds composed of 6-membered aromatic rings containing carbons and nitrogens (e.g.“c1ccccc1”, “c1cncncnc1”, “c1ccncc1”, “c1ncncn1”) are common, constituting 22% (67 out of 296) of these molecules.
However, Bemis-Murcko scaffolds are defined only for connected compounds with ring systems, and 186 molecules (almost 20% of the ApisTox dataset) violate this condition (141 have multiple fragments, 45 do not have any rings). In particular, it means that the ones with disconnected components could share an entire subgraph with those with scaffolds available, and the scaffold split method would not be able to detect that.
Next, we check the distribution of the functional groups, also called fragments in RDKit. In Figure 6, we visualize ten most discriminative functional groups, i.e. those with the largest frequency difference between pesticides toxic and non-toxic to honey bees. We see that there is a high proportion of insecticide-specific functional groups in toxic molecules. Phosphorus is very commonly found in instecticides, most often in a form of organophosphates [63, 64] (phos_ester and phos_acid fragments). Rich presence of sulfide fragment is explained by common usage of sulphur in pesticides, in particular fungicides [65, 66]. Pyrethrins and pyrethroids constitute a large group of insecticides [67], to which bees are very sensitive [68], containing an ester group and also very often a cyano group (ester and nitrile fragments). Neonicotinoids are yet another large group of insecticides, very rich in nitrogen, within which a hydrazine group is often detected, as well as a guanidinium group [69, 70] (hdrzine and guanido fragments). All those results show that ApisTox data aligns with literature on pesticide toxicity of honey bees.
@@
Lastly, we apply molecular frequent subgraph mining, in order to identify discriminative substructures, i.e. both frequent overall and at the same time much more frequent among toxic molecules than non-toxic ones (or vice versa). This is a more data-driven solution compared to functional groups, since we derive subgraphs from the data itself. It can also detect smaller subgraphs compared to typical functional groups. Again, we expect the data to reflect the usage of common chemical elements and their influence on pesticide toxicity to honey bees.
We use the MoSS tool [71] with default settings to mine the most common subgraphs. We then identify their frequency in toxic and non-toxic classes, and select ten with the largest frequency difference between classes. Results are presented in Figure 7. The results again align with literature, with phosphate-containing subgraphs much more common among toxic pesticides, and sulfur often found in both types of compounds.
@@
The important application of ApisTox dataset is understanding and predicting agrochemical compounds’ toxicity for honey bees, using data mining and machine learning (ML) methods. Those tasks can be divided into unsupervised and supervised applications, depending on whether they explicitly model the dependence between molecule features and selected targets or not.
@@
Analytical tasks include detecting functional groups, scaffolds, binding sites, and other molecular fragments that influence the toxicity of chemicals for honey bees. Those analyses can be carried out either on the entirety of the data, or on its subsets, e.g. for herbicides, fungicides, or insecticides only. Researching those factors can lead to better understanding of chemical space of pesticides non-toxic for honey bees, and therefore to design of safer pesticides in the future. Techniques like graph clustering, dimensionality reduction, and frequent subgraph mining can be utilized [72, 73]. We stress that those methods, even when used as unsupervised learning techniques, often have hyperparameters and settings to tune, and they should also be validated on external data not used during the initial analysis. Therefore, the provided train-test splits still have to be used, even if toxicity labels are not utilized. In this regard, time split would be especially useful, as a direct approximation of the process of developing new agrochemicals.
@@
Predictive algorithms will most likely model the relation between structure of molecules (SMILES column) and either label or ppdb_level column as the target. They can be useful tools for pesticide design, acting as ML-based filters for new agrochemical candidates. Substances with high predicted probability of honey bee toxicity can be ruled out early, lowering costs and accelerating the process. Such models should utilize only the training data for parameter estimation. For hyperparameter tuning, cross-validation with stratified sampling is recommended, due to class imbalance and relatively small dataset size. Test data must not be used at any point until the final validation of generalization performance. We recommend using MaxMin split for this purpose, since it covers the chemical space of the dataset more uniformly than other splits, requiring good generalization across the entire domain of the dataset [48, 49].
@@
For measuring the performance of predictive models, at least two or three metrics should be used, taking into consideration class imbalance and reporting different aspects of model performance. In particular, we recommend Area Under Receiver Operating Characteristic curve (AUROC), since it works well with imbalanced data and also takes into consideration probabilistic outputs of classifiers, as well as Matthews Correlation Coefficient (MCC), which can sometimes detect model failures despite high AUROC value [74, 75]. Among other popular metrics, F1-score, precision, and recall can be used.
@@
ApisTox can also be used as a benchmark dataset for molecular graph classification. Since datasets and benchmarks in this area come almost exclusively from medicinal chemistry, performance of many models, like molecular fingerprints, graph kernels, and graph neural networks (GNNs), has been evaluated exclusively on tasks directly related to pharmacological drug design. Our dataset allows validation of the generalization performance of such models on new domains of agrochemistry and ecotoxicology. In this context, usage of unified and predefined split and metrics is of paramount importance, to allow comparison of different models. We recommend usage of MaxMin split, as well as reporting both AUROC and MCC.
@@
When using ppdb_level as target variable, we note that this is not a three class classification problem, but rather ordinal regression, also known as ordinal classification [76]. Toxicity levels are ordered integers, and classes 0 and 2 (non-toxic and highly toxic) are more distant than 1 and 2 (moderately toxic and highly toxic). Therefore, appropriate models should be used, e.g. ordinal logit model instead of logistic regression. In this case, additionally reporting regression metrics like MAE (Mean Absolute Error) and RMSE (Root Mean Squared Error) is recommended, with additional corrections for class imbalance [77].
@@
Code is available on GitHub at https://github.com/j-adamczyk/ApisTox_dataset. Code uses Python 3.10. To ensure full reproducibility, we pinned all external library dependencies (including transitive dependencies) using Poetry dependency manager [78]. We also include poetry.lock file with all dependency versions, as well as requirements.txt file exported from it.
@@ Acknowledgements
Research project supported by program “Excellence initiative – research university” for the AGH University of Krakow.
@@
J.A. and J.P.: Conceptualization, Investigation, Methodology, Data Curation, Software, Validation, Visualization, Writing - Original Draft, Writing - Review & Editing. P.S.: Conceptualization, Supervision, Methodology, Writing - Original Draft, Writing - Review & Editing. All authors reviewed and approved the manuscript.
@@
ApisTox - supplementary information
Herbicides Fungicides Insecticides Other agrochemicals Unknown
Split type Train Test Train Test Train Test Train Test Train Test
Stratified random 35% 32% 18% 22% 22% 22% 21% 21% 6% 4%
Time 32% 41% 19% 21% 24% 15% 22% 16% 5% 7%
MaxMin 37% 21% 18% 21% 22% 20% 18% 33% 5% 6%
Table 6: Pesticide type distributions for different splits.
Note that pesticides can have more than one type, so it is possible for a given split type for train or test to sum up to more than 100%.